OpenStudy (anonymous):

R=3K - Q, for k this one is hard and i already have the work i just want to know if im right

jimthompson5910 (jim_thompson5910):

R=3K - Q R+Q=3K - Q+Q R + Q = 3K 3K = R + Q 3K/3 = (R + Q)/3 K = (R + Q)/3 So the solution is K = (R + Q)/3